# On the contraction properties of Sinkhorn semigroups

### v2

@@ D.2 Some quantitative estimates
When $\jmath({\cal Q})>0$, for any $n\geq 0$ we have the uniform estimates
$\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\cal Q})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\cal Q})^{-2}.$ (127)
@@
Recalling that $\jmath({\cal S}_{n})\wedge\jmath({\cal Q})\geq\hbar({\cal S}_{n})=\hbar({\cal Q})$, for any given $x\in\mathbb{X}$ we have
$\hbar({\cal Q})~{\cal S}_{2n}(x,dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})=({\cal S}_{2n-1}{\cal S}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})}~{\cal S}_{2n}(x,dy_{2}).$
@@
$\hbar({\cal Q})^{2}~{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\cal S}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~{\cal S}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
@@
$\hbar({\cal Q})^{2}~\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~\nu_{V}(dy_{2}).$
@@
$\hbar({\cal Q})^{2}~\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\cal Q})^{2}}~\pi_{2n}(dy_{2}).$
@@
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\cal Q})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\cal Q})^{-2}.$
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
$\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$ (128)
@@
Using (125) and (53), we see that, for any $n\geq n_{0}$,
$\frac{d\pi_{2n}}{d\nu_{V}}$ $\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~\left({\cal S}^{\circ}_{2n}\ldots{\cal S}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
@@
In addition, when Condition ${\cal D}_{\varphi,\psi}({\cal S})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
@@
If Condition ${\cal D}_{\varphi,\psi}({\cal S})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.8, the drift condition (53) also holds for any $n>n_{0}$, which implies that
@@
Arguing as in the proof of Proposition D.2, now combining the drift conditions (60) with (126), we find that
@@
Under the stronger Condition ${\cal H}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
${\cal W}_{V,U}(x):=\log{{\cal Q}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\cal W}^{U,V}(y):=\log{{\cal R}_{U}(e^{W^{V}})(y)},$ (129)
with the integrated cost functions $(W_{U},W^{V})$ defined in (28) and the integral operators
${\cal Q}_{V}(x,dy)=e^{-W(x,y)}~\nu_{V}(dy)\quad\mbox{\rm and}\quad{\cal R}_{U}(y,dx)=e^{-W(x,y)}~\lambda_{U}(dx).$
Next technical lemma is pivotal.
@@
$\begin{array}[]{rcl}\log{{\cal Q}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\cal W}^{U,V}\geq\log{{\cal R}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$ (130)
@@
$\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\cal W}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\cal W}^{U,V}-\nu_{V}(V).\end{array}$ (132)
@@
Formula (21) yields the following monotone properties
$\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~|~\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~|~\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$ (133)
@@
By (133) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
$\log{{\cal R}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~\lambda_{U}(dx)~\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
@@
using Jensen’s inequality again. This ends the proof of (130).
The integral equations in (18) imply that for any $n\geq 0$ we have
@@
Using again (18) we see that
$U_{2(n+1)}-U=\log{{\cal Q}(e^{-V_{2(n+1)}})}$ $\leq$ $\log{{\cal Q}_{V}(e^{W_{U}})}={\cal W}_{V,U}\quad\text{and}$
$V_{2(n+1)}-V=\log{{\cal R}(e^{-U_{2n}})}$ $\leq$ $-\nu_{V}(V)+\log{{\cal R}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\cal W}^{U,V}.$
@@
If Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}\vee\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$ (134)
@@
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~e^{W^{V}}.$ (135)
@@
$\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\cal Q}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
@@
$\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\cal R}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\cal H}_{\delta}(U,V)$. This ends the proof of (134).
Using (132) we see that, for any $n\geq 1$,
${\cal W}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\cal W}_{V,U}+W^{V}-\nu_{V}(V)$
@@
${\cal W}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\cal W}^{U,V}+W_{U}-\nu_{V}(V).$
Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (140) and (141) in the proof of the commutation formulae (123), we readily obtain the estimates
$e^{\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W_{U}}$
@@
$e^{\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\cal W}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W^{V}}$
@@
The functions $({\cal W}_{V,U},{\cal W}^{U,V})$ defined in (129) can be computed explicitly for the the Gaussian model defined in (87) and (88).
@@
${\cal W}_{V,U}(x)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~(\beta_{\tau}~(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
@@
${\cal W}^{U,V}(y)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$ (136)
@@
Using (135), Condition ${\cal H}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
$e^{-\delta V}~\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~e^{-{\cal V}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~~\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~e^{-{\cal U}_{\delta}}$
@@
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\cal B}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\cal B}_{\varphi}(\mathbb{X}).$

### v3

@@ D.2. Some quantitative estimates
When $\jmath({\mathcal{Q}})>0$, for any $n\geq 0$ we have the uniform estimates
(D.5) $\|\frac{d\pi_{2n}}{d\nu_{V}}\|\vee\|\frac{d\nu_{V}}{d\pi_{2n}}\|\leq\hbar({\mathcal{Q}})^{-2}\quad\mbox{and}\quad\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|\vee\|\frac{d\lambda_{U}}{d\pi_{2n+1}}\|\leq\hbar({\mathcal{Q}})^{-2}.$
@@
Recalling that $\jmath({\mathcal{S}}_{n})\wedge\jmath({\mathcal{Q}})\geq\hbar({\mathcal{S}}_{n})=\hbar({\mathcal{Q}})$, for any given $x\in\mathbb{X}$ we have
$\hbar({\mathcal{Q}})~{\mathcal{S}}_{2n}(x,dy_{2})\leq{\mathcal{S}}_{2n}^{\circ}(y_{1},dy_{2})=({\mathcal{S}}_{2n-1}{\mathcal{S}}_{2n})(y_{1},dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})}~{\mathcal{S}}_{2n}(x,dy_{2}).$
@@
$\hbar({\mathcal{Q}})^{2}~{\mathcal{S}}_{2n}^{\circ}(\overline{y}_{1},dy_{2})\leq{\mathcal{S}}_{2n}^{\circ}(y_{1},dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~{\mathcal{S}}_{2n}^{\circ}(\overline{y}_{1},dy_{2})$
@@
$\hbar({\mathcal{Q}})^{2}~\nu_{V}(dy_{2})\leq\pi_{2n}(dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~\nu_{V}(dy_{2}).$
@@
$\hbar({\mathcal{Q}})^{2}~\pi_{2n}(dy_{2})\leq\nu_{V}(dy_{2})\leq\frac{1}{\hbar({\mathcal{Q}})^{2}}~\pi_{2n}(dy_{2}).$
@@
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq\hbar({\mathcal{Q}})^{-2}\quad\text{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq\hbar({\mathcal{Q}})^{-2}.$
@@
If Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ holds for some $n_{0}\geq 1$ then there exists $c_{\epsilon}>0$ such that
(D.6) $\sup_{n\geq n_{0}}\|\frac{d\pi_{2n}}{d\nu_{V}}\|_{\psi}\leq c_{\epsilon}~\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}\quad\mbox{\rm and}\quad\sup_{n\geq n_{0}}\|\frac{d\pi_{2n+1}}{d\lambda_{U}}\|_{\varphi}\leq c_{\epsilon}~\|\frac{d\pi_{2n_{0}-1}}{d\lambda_{U}}\|_{\varphi}.$
@@
Using (D.3) and (4.6), we see that, for any $n\geq n_{0}$,
$\frac{d\pi_{2n}}{d\nu_{V}}$ $\leq$ $\|\frac{d\pi_{2(n_{0}-1)}}{d\nu_{V}}\|_{\psi}~\left({\mathcal{S}}^{\circ}_{2n}\ldots{\mathcal{S}}^{\circ}_{2n_{0}}\right)\left(\psi\right)$
@@
In addition, when Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ holds for some $n_{0}\geq 0$, there exists $c_{\epsilon}>0$ such that
@@
If Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}})$ is satisfied for some integer $n_{0}\geq 0$ then, by Theorem 4.6, the drift condition (4.6) also holds for any $n>n_{0}$, which implies that
@@
Arguing as in the proof of Proposition D.2, now combining the drift conditions (4.13) with (D.4), we find that
@@
Under the stronger Condition ${\mathcal{H}}_{\delta}(U,V)$, more refined estimates can be obtained. With that end, consider the functions
(D.7) ${\mathcal{W}}_{V,U}(x):=\log{{\mathcal{Q}}_{V}(e^{W_{U}})(x)}\quad\mbox{\rm and}\quad{\mathcal{W}}^{U,V}(y):=\log{{\mathcal{R}}_{U}(e^{W^{V}})(y)},$
with the integrated cost functions $(W_{U},W^{V})$ defined in (2.14) and the integral operators
${\mathcal{Q}}_{V}(x,dy)=e^{-W(x,y)}~\nu_{V}(dy)\quad\mbox{\rm and}\quad{\mathcal{R}}_{U}(y,dx)=e^{-W(x,y)}~\lambda_{U}(dx).$
The next technical lemma is pivotal.
@@
(D.8) $\begin{array}[]{rcl}\log{{\mathcal{Q}}(\exp{(-V_{2n})})}&\geq&-W^{V}+\nu_{V}(V)\quad\text{and}\\ &&\\ -\nu_{V}(V)+{\mathcal{W}}^{U,V}\geq\log{{\mathcal{R}}(\exp{(-U_{2n})})}&\geq&-W_{U},\end{array}$
@@
(D.10) $\begin{array}[]{rcccl}U-W^{V}+\nu_{V}(V)&\leq&U_{2n}&\leq&U+{\mathcal{W}}_{V,U}\quad\text{and}\\ &&&&\\ V-W_{U}&\leq&V_{2n}&\leq&V+{\mathcal{W}}^{U,V}-\nu_{V}(V).\end{array}$
@@
Formula (2.7) yields the following monotone properties
(D.11) $\begin{array}[]{rcccccl}\nu_{V}(V_{2(n+1)})&\leq&\nu_{V}(V_{2n})&=&\nu_{V}(V_{2(n+1)})+\mbox{\rm Ent}(\nu_{V}~|~\pi_{2n})&\leq&\nu_{V}(V_{0})=0,\\ &&&&&&\\ \lambda_{U}(U_{2(n+1)})&\leq&\lambda_{U}(U_{2n})&=&\lambda_{U}(U_{2(n+1)})+\mbox{\rm Ent}(\lambda_{U}~|~\pi_{2n+1})&\leq&\lambda_{U}(U_{0}).\end{array}$
@@
By (D.11) we also have $\lambda_{U}(U)\geq\lambda_{U}(U_{2n})$ and, therefore, for any $n\geq 0$ we obtain
$\log{{\mathcal{R}}(\exp{(-U_{2n})})(y)}$ $=$ $\log{\int~\lambda_{U}(dx)~\exp{(-W(x,y)+U(x)-U_{2n}(x))}}$
@@
using Jensen’s inequality again. This ends the proof of (D.8).
The integral equations in (2.4) imply that for any $n\geq 0$ we have
@@
Using again (2.4) we see that
$U_{2(n+1)}-U=\log{{\mathcal{Q}}(e^{-V_{2(n+1)}})}$ $\leq$ $\log{{\mathcal{Q}}_{V}(e^{W_{U}})}={\mathcal{W}}_{V,U}\quad\text{and}$
$V_{2(n+1)}-V=\log{{\mathcal{R}}(e^{-U_{2n}})}$ $\leq$ $-\nu_{V}(V)+\log{{\mathcal{R}}_{U}(e^{W^{V}})}=-\nu_{V}(V)+{\mathcal{W}}^{U,V}.$
@@
If Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1]$, then
(D.12) $\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}\vee\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}<\infty.$
@@
(D.13) $\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\leq c~e^{W_{U}}\quad\mbox{and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\leq c~e^{W^{V}}.$
@@
$\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}$ $=$ $\|{\mathcal{Q}}_{V}(e^{W_{U}})\|_{\mathbb{L}_{\infty}(\lambda)}$
@@
$\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\nu)}=\|{\mathcal{R}}_{U}(e^{W^{V}})\|\leq\|e^{-\delta U}\|_{\mathbb{L}_{\infty}(\lambda)}~\|e^{-W}\|_{\mathbb{L}_{\infty}(\lambda\otimes\nu)}~\lambda_{(1-\delta)U}\left(e^{W^{V}}\right),$
where $\lambda_{(1-\delta)U}\left(e^{W^{V}}\right)\vee\nu_{(1-\delta)V}\left(e^{W_{U}}\right)<\infty$ by Condition ${\mathcal{H}}_{\delta}(U,V)$. This ends the proof of (D.12).
Using (D.10) we see that, for any $n\geq 1$,
${\mathcal{W}}_{V,U}-W^{V}+\nu_{V}(V)\leq U_{2(n+1)}-U_{2n}\leq{\mathcal{W}}_{V,U}+W^{V}-\nu_{V}(V)$
@@
${\mathcal{W}}^{U,V}-W_{U}+\nu_{V}(V)\leq V_{2(n+1)}-V_{2n}\leq{\mathcal{W}}^{U,V}+W_{U}-\nu_{V}(V).$
Then, recalling that $U_{2n+1}=U_{2n}$ and $V_{2(n+1)}=V_{2n+1}$ and using Eqs. (E.1) and (E.2) in the proof of the commutation formulae (D.1), we readily obtain the estimates
$e^{\nu_{V}(V)}\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W_{U}}\leq\frac{d\pi_{2n}}{d\nu_{V}}\leq e^{-\nu_{V}(V)}\|e^{{\mathcal{W}}^{U,V}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W_{U}}$
@@
$e^{\nu_{V}(V)}\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}^{-1}~e^{-W^{V}}\leq\frac{d\pi_{2n+1}}{d\lambda_{U}}\leq e^{-\nu_{V}(V)}\|e^{{\mathcal{W}}_{V,U}}\|_{\mathbb{L}_{\infty}(\lambda)}~e^{W^{V}}$
@@
The functions $({\mathcal{W}}_{V,U},{\mathcal{W}}^{U,V})$ defined in (D.7) can be computed explicitly for the the Gaussian model defined in (5.6) and (5.7).
@@
${\mathcal{W}}_{V,U}(x)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\sigma_{\beta})+\frac{1}{2}~(\beta_{\tau}(x-m))^{\prime}(\overline{\sigma}-\tau)~(\beta_{\tau}~(x-m))+(\overline{m}-m_{0})^{\prime}\beta_{\tau}(x-m)$
@@
(D.14) ${\mathcal{W}}^{U,V}(y)=\frac{1}{2}~\mbox{\rm Tr}(\tau^{-1}\overline{\sigma})+\frac{1}{2}(y-\overline{m})^{\prime}\left(\beta_{\tau}\,\sigma\,\beta_{\tau}^{\prime}-\tau^{-1}\right)(y-\overline{m})-(y-\overline{m})^{\prime}\tau^{-1}(\overline{m}-m_{0}).$
@@
Using (D.13), Condition ${\mathcal{H}}_{\delta}(U,V)$ ensures that for any $n\geq 1$ we have
$e^{-\delta V}~\left(\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\right)\leq c~e^{-{\mathcal{V}}_{\delta}}\quad\mbox{\rm and}\quad e^{-\delta U}~~\left(\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\right)\leq c~e^{-{\mathcal{U}}_{\delta}}$
@@
$\frac{d\pi_{2n}}{d\nu_{V}}\vee\frac{d\nu_{V}}{d\pi_{2n}}\in{\mathcal{B}}_{\psi}(\mathbb{Y})\quad\mbox{\rm and}\quad\frac{d\pi_{2n+1}}{d\lambda_{U}}\vee\frac{d\lambda_{U}}{d\pi_{2n+1}}\in{\mathcal{B}}_{\varphi}(\mathbb{X}).$